# Magnon thermal Hall effect in collinear ferrimagnets

### v3

@@ Abstract
In this paper we theoretically discuss thermal Hall effect of magnons in insulating Néel ordered antiferromagnets at zero external magnetic field. We show that for compensated Néel order the non-zero thermal Hall effect will occur inthe absence of any symmetry between the two magnetic sublattices, thus making the system ferrimagnetic. The thermal Hall effect of magnons will be non-zero by a virtue of the spin-momentum splitting of the magnon spectrum due to the Dzyaloshinskii-Moriya interaction as well as anisotropic second-nearest exchange interaction different in the two magnetic sublattices, both corresponding to the broken symmetry. We construct a theoretical model in which an external electric field may change the symmetry of the antiferromagnetic system thus altering the thermal Hall effect of magnons.
@@
There are two types of antiferromagnets in the classification scheme made by Turov Turov1965 ; Turov1990 of antiferromagnets with equal (connected by a symmetry operation) magnetic sublattices. Namely, these are genuine antiferromagnets with zero net magnetization and Dzyaloshinskii’s weak ferromagnets Dzyaloshinskii1958 with small non-zero magnetic moment. In Dzyaloshinskii’s weak ferromagnets crystal symmetry and special directions of the Néel vector allow for the magnetic moment. In addition, there are antiferromagnetic ferrimagnets in which case magnetic sublattices with equal in magnitude and antialigned spins aren’t connected by any symmetry operation.
@@
In addition, more insight comes from the point of view of symmetries of the magnetic unit cell of an antiferromagnet. For example, the local non-magnetic environment of the Néel order in metallic antiferromagnets may give rise to spin splitting of conducting fermions that interact with the Néel order Noda2016 ; Okugawa2018 ; HayamiYanagiKusunose2019 ; AHE_AFM ; Ahn2019 ; Rashba2020 . The spin-splitting may be antiferromagnetic or ferromagnetic depending on what symmetries the non-magnetic environment breaks. If the non-magnetic environment is such that the magnetic sublattices are connected by a combination of time-reversal and translation, then the conducting fermions are spin degenerate with zero net magnetic moment. If the non-magnetic environment is such that only a combination of certain rotations or mirror operations and time-reversal connects the magnetic sublattices, then the spin-splitting of conducting fermions will be of the antiferromagnetic type corresponding to the rotation or mirror symmetry. Such spin-splittings might be of the $d$-,$g$-,$i$- types Noda2016 ; Okugawa2018 ; HayamiYanagiKusunose2019 ; AHE_AFM ; Ahn2019 ; Rashba2020 ; HayamiYanagiKusunose2020 or mirror-symmetric Zyuzin2025 . It was shown that the spin-splitting together with spin-orbit coupling may result in the anomalous Hall effect TurovShavrov ; Turov1990 ; AHE_AFM ; Naka2019 ; Naka2020 in antiferromagnets. Such antiferromagnets are Dzyaloshinskii’s weak ferromagnets in which the conducting fermions are responsible for magnetization while the Néel order is intact.
@@
In the case when the local non-magnetic environment breaks all symmetries between the magnetic sublattices, the spin-splitting of conducting fermions will acquire a ferromagnetic contribution which would result in finite spin-polarization of conducting fermions Zyuzin2024 ; Zyuzin2025 . Such antiferromagnets are called ferrimagnets. In a ferrimagnet studied in Zyuzin2025 the spin-splitting of conducting fermions was a sum of antiferromagnetic mirror-symmetric and ferromagnetic terms. It was shown that ferrimagnets have an anomalous Hall effect Zyuzin2025 .
@@
It is then natural to understand implications of the local non-magnetic environment of the Néel order in insulating antiferromagnets on the thermal Hall effect (THE) of magnons MatsumotoShindouMurakamiPRB2014 ; LaurellFiete ; Mook2019 ; ZyuzinKovalevPRL2016 ; Pires ; Mook in the absence of external magnetic field. In ZyuzinKovalevPRL2016 the THE of magnons was shown to be absent in the collinear antiferromagnet on a honeycomb lattice in which case the two sublattices are connected by a combination of time-reversal and $\frac{\pi}{3}$ rotation. In Pires antiferromagnetic magnons have been studied in collinear antiferromagnet on a checkerboard lattice in which case the two sublattices are connected by a combination of time-reversal and $\frac{\pi}{2}$ rotation. The THE of magnons is absent in this system as well because of the symmetry between the magnetic sublattices. In Mook a three-dimensional antiferromagnet on a rutile lattice was studied and it was found that non-zero THE of magnons is due to first-nearest neighbor Dzyaloshinskii-Moriya interaction. It does appear that the system studied in Mook belongs to the Dzyaloshinskii’s weak ferromagnet type Dzyaloshinskii1958 , hence the obtained non-zero magnon THE.
@@
In this paper we wish to understand the mechanism of magnon THE in insulating ferrimagnets. For that we studied theoretical two-dimensional model of an antiferromagnet proposed in Zyuzin2025 . We show that finite magnon THE can be found in insulating collinear antiferromagnets in which there is no symmetry between the two magnetic sublattices, i.e. in ferrimagnets. Absence of symmetry is due to the non-symmetric environment of the non-magnetic atoms. In addition to this, the first-nearest neighbor Dzyaloshinskii-Moriya interaction (DMI) consistent with the asymmetry of the magnetic sublattices as well as the second-nearest Heisenberg exchange interaction (HEI) are required for the non-zero magnon THE. Furthermore, we suggest that the external electric field can vary the symmetry of the lattice when the unit cell of the insulating antiferromagnet can have a finite electric dipole moment. The system can then be driven into a low symmetry phase by the electric field. In that way, by varying the direction of the field, the magnon THE can be altered. This can serve as a powerful experimental tool to probe the structure of antiferromagnetic magnons.
@@
Figure 1: The models of two antiferromagnetic systems are depicted. The red and blue sites represent the spin-up and spin-down magnetic sublattices of the Néel order. The systems include a nearest-neighbor Heisenberg exchange interaction (HEI, $J>0$) and Dzyaloshinskii-Moriya interaction (DMI). Further-neighbor HEI and DMI are indicated by dashed lines. In a simplifying assumption, we exclude any interactions across the green atom. The arrows on the bonds denote the sign of the DMI. The left model possesses a symmetry connecting the two magnetic sublattices, whereas this symmetry is absent in the right model.
@@
We study two insulating antiferromagnetic systems with the Néel order on the honeycomb-like lattices shown in Fig. (1). These models were studied in Zyuzin2025 for the situation when the systems are metallic. The red and blue sites correspond to ${\bf S}_{1/2}=\pm{\bf S}$ magnetic sublattices. The green atom is non-magnetic, and its role is to alter the first-nearest neighbor Heisenberg (HEI) and Dzyaloshinskii-Moriya (DMI) interactions in its vicinity through the spin-orbit interaction. We also assume that there is no second-nearest neighbor HEI or DMI through the green atom. This is a simplification; however, our results below remain unchanged even if these interactions are non-zero, provided their magnitude differs from that of the corresponding interactions on bonds without the green atom. Exchange Hamiltonian of an antiferromagnet model shown in Fig. (1) is
$H$ $=J\sum_{\langle ij\rangle}S_{i}^{z}S_{j}^{z}+J\cos(\theta)\sum_{\langle ij\rangle}\left(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}\right)$
$+J\sin(\theta)\sum_{\langle ij\rangle}\nu_{ij}\left(S_{i}^{x}S_{j}^{y}-S_{i}^{y}S_{j}^{x}\right)$
@@
where $J>0$ and $D\equiv J\sin(\theta)$ is the HEI and DMI between the first neighbors labeled by the $\langle..\rangle$ in the sum. In Fig. (1) $\nu_{ij}=\pm 1$ for the bond along the arrow. If the arrow on the $\langle ij\rangle$ bond is reversed, then $\nu_{ij}=-\nu_{ji}$. If there is no arrow on the bond then $\nu_{ij}=0$. The signs are chosen in correspondence with the underlying spin-orbit coupling. In addition we have added $J^{\prime}$ and $D^{\prime}$, which is the HEI and DMI correspondingly between the second-nearest neighbors, labeled by the $\langle\langle..\rangle\rangle$. We assume that $J^{\prime}<0$, i.e. it is of the ferromagnetic sign. Such a sign supports the Néel order. In addition to this, second-nearest neighbor interactions may be bond and sublattice dependent. The details of these interactions will be specified below for two different models shown in Fig. (1). We stress that there is no external magnetic field in the system.
@@
The thermal Hall effect of antiferromagnetic magnons has been derived in MatsumotoShindouMurakamiPRB2014 ( also see comment ). Magnong thermal Hall effect is defined as the heat current Luttinger1964 carried by the magnons, which flows in the transverse direction to the applied temperature gradient $\nabla_{\beta}\chi$, namely,
$j^{\mathrm{TH}}_{\alpha}=\frac{1}{2TV}\sum_{{\bf k};n}(\hat{\Omega}_{{\bf k};\alpha\beta})_{nn}c_{2}\left[(\hat{\sigma}_{3}\hat{E}_{\bf k})_{nn}\right]\nabla_{\beta}\chi,$ (16)
@@
$j^{\mathrm{TH}}_{\alpha}=\frac{1}{TV}\sum_{{\bf k};n}\Omega_{{\bf k};\alpha\beta}\left[c_{2}\left(\delta_{\bf k}+\varepsilon_{\bf k}\right)+c_{2}\left(-\delta_{\bf k}+\varepsilon_{\bf k}\right)\right]\nabla_{\beta}\chi.$ (20)
From ZyuzinKovalevPRL2016 it is known that odd in momentum Berry curvature of antiferromagnetic magnons may be due to the lattice structure (honeycomb lattice), while from Pires (checkerboard lattice with d-wave DMI and finite $\delta_{\bf k}$) it is known that even in momentum Berry curvature may be due to the even in momentum DMI between the sublattices. However, both models, ZyuzinKovalevPRL2016 and Pires don’t allow for non-zero magnon thermal Hall effect because of the symmetry between the magnetic sublattices. Then, in Mook an antiferromagent on the rutile lattice was shown to have finite THE due to the even in momentum Berry curvature, similarly to Pires , originating from DMI. It appears that the antiferromagnet on a ruile lattice is the Dzyaloshinskii’s weak ferromagnet Dzyaloshinskii1958 . Here we show that there is another possibility for non-zero magnon thermal Hall in antiferromagnets shown in Fig. (1) with odd in momentum magnon Berry curvature due to nearest neighbor HEI, and with the even in momentum Berry curvature due to the nearest neighbor DMI. Below we will show that magnon THE will be non-zero only when all the symmetries between the two magnetic sublattices are broken, and hence the system is a ferrimagnet in this case. Let us discuss the details of models I and II.
@@
It is clear that the integration over the momentum will make the THE to vanish in this model. It is instructive to see how the Dzyaloshinskii-Turov symmetry Dzyaloshinskii1958 ; Turov1965 arguments support our result. We introduce a Néel vector ${\bf L}={\bf S}_{1}-{\bf S}_{2}$ and the magnetization ${\bf M}={\bf S}_{1}+{\bf S}_{2}$. We then analyze whether a Dzyaloshinskii’s invariant of the $M_{\alpha}L_{\beta}$ (where $\alpha,\beta=x,y,z$) form can be realized in the system. For it to be non-zero, it has to be invariant under all symmetries of the system. If it is non-zero, it would mean that $L_{\alpha}$ can generate a finite magnetic moment $M_{\beta}$ in the system. In two dimensions we are expecting only $M_{z}$ to be generated. Both vectors transform as pseudovectors, and in addition, $\bf L$} changes sign if the magnetic sublattices are exchanged. For $L_{y}$ and $L_{z}$, a reflection in the yzy-z plane which crosses the vertical bond in the center is the symmetry connecting the two magnetic sublattices. Under this symmetry operation there is no Dzyaloshinskii’s invariant of the $M_{z}L_{z}$ or $M_{z}L_{y}$ form. Indeed, $L_{z}$ and $L_{y}$ are invariant, while $M_{z}$ changes sign. Furthermore, a symmetry of $\pi$ rotation about the center of the hexagon and a reflection in the xzx-z plane which crosses the center of the hexagon eliminates $M_{z}L_{x}$ from the list of Dzyaloshinskii’s invariants. Indeed, $L_{x}$ doesn’t change the sign while $M_{z}$ does under this symmetry operation. In addition, it is worth noticing that a reflection in the xyx-y plane and time-reversal operation elimintaes the $M_{z}L_{x/y}$ combinations as well. All in all, the system in the left of Fig. (1) is a genuine antiferromagnet with zero magnetic moment, and magnon THE is not expected in this system.

### v4

# Magnon thermal Hall effect in collinear antiferromagnets

@@ Abstract
In this paper we theoretically discuss thermal Hall effect of magnons in insulating Néel ordered antiferromagnets at zero external magnetic field. We show that for compensated Néel order the non-zero thermal Hall effect will occur in the absence of any symmetry between the two magnetic sublattices, thus making the system ferrimagnetic. We then show that collinear Dzyaloshinskii’s weak ferromagnets, in which there is a symmetry connecting the magnetic sublattices, also show magnon thermal Hall effect. The thermal Hall effect of magnons will be non-zero by a virtue of the spin-momentum splitting of the magnon spectrum due to the Dzyaloshinskii-Moriya interaction as well as second-nearest exchange interaction different in the two magnetic sublattices, both corresponding to the broken symmetries that lead to the Dzyaloshinskii’s invariant. We construct a theoretical model in which an external electric field may change the symmetry of the antiferromagnetic system thus altering the thermal Hall effect of magnons.
I Introduction
@@
There are two types of antiferromagnets in the classification scheme made by Turov Turov1965 ; Turov1990 of antiferromagnets with equal (connected by a symmetry operation) magnetic sublattices. Namely, these are genuine antiferromagnets with zero net magnetization and Dzyaloshinskii’s weak ferromagnets BorovikRomanov1957 ; Dzyaloshinskii1958 with small non-zero magnetic moment. In Dzyaloshinskii’s weak ferromagnets crystal symmetry and special directions of the Néel vector allow for the magnetic moment. In addition, there are antiferromagnetic ferrimagnets in which case magnetic sublattices with equal in magnitude and antialigned spins aren’t connected by any symmetry operation.
@@
In addition, more insight comes from the point of view of symmetries of the magnetic unit cell of an antiferromagnet. For example, the local non-magnetic environment of the Néel order in metallic antiferromagnets may give rise to spin splitting of conducting fermions that interact with the Néel order Noda2016 ; Okugawa2018 ; HayamiYanagiKusunose2019 ; Naka2019 ; AHE_AFM ; Ahn2019 ; Rashba2020 ; HayamiYanagiKusunose2020 ; Zyuzin2025 . The spin-splitting may be antiferromagnetic or ferromagnetic depending on what symmetries the non-magnetic environment breaks. If the non-magnetic environment is such that the magnetic sublattices are connected by a combination of time-reversal and translation or inversion, then the conducting fermions are spin degenerate with zero net magnetic moment. If the non-magnetic environment is such that only a combination of certain rotations or mirror operations and time-reversal connects the magnetic sublattices, then the spin-splitting of conducting fermions will be of the antiferromagnetic type corresponding to the rotation or mirror symmetry. Such spin-splittings might be of the $d$-,$g$-,$i$- types Noda2016 ; Okugawa2018 ; HayamiYanagiKusunose2019 ; Naka2019 ; AHE_AFM ; Ahn2019 ; Rashba2020 ; HayamiYanagiKusunose2020 or mirror-symmetric Zyuzin2025 . The d-wave spin splitting has been experimentally observed in Ref. KharchenkoBibikEremenko1985 ; EremenkoKharchenko1987 through the d-wave Hall effect VorobevZyuzin2024 . It was shown that the spin-splitting together with spin-orbit coupling may result in the anomalous Hall effect TurovShavrov ; Turov1990 ; AHE_AFM in antiferromagnets. Such antiferromagnets are Dzyaloshinskii’s weak ferromagnets in which the conducting fermions are responsible for magnetization while the Néel order is intact.
@@
In the case when the local non-magnetic environment breaks all symmetries between the magnetic sublattices, the spin-splitting of conducting fermions will acquire a ferromagnetic contribution which would result in finite spin-polarization of conducting fermions Naka2020 ; Zyuzin2024 ; Zyuzin2025 . Such antiferromagnets are called ferrimagnets. In a ferrimagnet studied in Zyuzin2025 the spin-splitting of conducting fermions was a sum of antiferromagnetic mirror-symmetric and ferromagnetic terms. It was shown that ferrimagnets have an anomalous Hall effect Zyuzin2025 .
@@
It is then natural to understand implications of the local non-magnetic environment of the Néel order in insulating antiferromagnets on the thermal Hall effect (THE) of magnons MatsumotoShindouMurakamiPRB2014 ; LaurellFiete ; Mook2019 ; ZyuzinKovalevPRL2016 ; Pires ; Mook in the absence of external magnetic field. In ZyuzinKovalevPRL2016 the THE of magnons was shown to be absent in the collinear antiferromagnet on a honeycomb lattice in which case the two sublattices are connected by a combination of time-reversal and $\frac{\pi}{3}$ rotation or by a combination of time-reversal and inversion. In Pires antiferromagnetic magnons have been studied in collinear antiferromagnet on a checkerboard lattice in which case the two sublattices are connected by a combination of time-reversal and $\frac{\pi}{2}$ rotation. The THE of magnons is absent in this system as well because of the symmetry between the magnetic sublattices. In Mook ; CommentAlter a three-dimensional antiferromagnet on a rutile lattice was studied and it was found that non-zero THE of magnons is due to first-nearest neighbor Dzyaloshinskii-Moriya interaction. The system studied in Mook ; CommentAlter belongs to the Dzyaloshinskii’s weak ferromagnet type Dzyaloshinskii1958 ; Turov1965 , hence the obtained non-zero magnon THE.
@@
In this paper we wish to systematically understand the mechanism of magnon THE in insulating collinear antiferromagnets. By collinearity we mean absence of any canting of the Néel order. We study theoretical two-dimensional model of an antiferromagnet proposed in Zyuzin2025 . This model allows to tune between different antiferromagnetic phases (genuine antiferromagnet, ferrimagnet and weak ferromagnet) in a simple way. We deduce Dzyaloshinskii’s invariants of existence of a finite magnetic moment in collinear antiferromagnets for ferrimagnet and weak ferromagnet models and show that the derived by us magnon THE in the models is consistent with the invariant. We first show that finite magnon THE can be found in insulating collinear antiferromagnets in which there is no symmetry between the two magnetic sublattices, i.e. in ferrimagnets. Absence of symmetry is due to the non-symmetric environment of the non-magnetic atoms. In addition to this, the first-nearest neighbor Dzyaloshinskii-Moriya interaction (DMI) consistent with the asymmetry of the magnetic sublattices as well as the second-nearest Heisenberg exchange interaction (HEI) different in the two magnetic sublattices are required for the non-zero magnon THE. Both ingredients are consistent with the Dzyaloshinskii’s invariant. We then show that magnon THE indeed appears in a weak ferromagnetic phase of the model proposed in Zyuzin2025 . Weak ferromagnet is a collinear antiferromagnet in which magnetic sublattices are connected to each other by some symmetry operation and which allow for the finite magnetic moment. The magnon THE in the weak ferromagnetic phase is also governed by DMI and second-nearest HEI consistent with the lattice symmetry and Dzyaloshinskii’s invariant.
Furthermore, we suggest that the external electric field can vary the symmetry of the lattice when the unit cell of the insulating antiferromagnet can have a finite electric dipole moment. The system can then be driven into a low symmetry phase by the electric field. In that way, by varying the direction of the field, the magnon THE can be altered. This can serve as a powerful experimental tool to probe the structure of antiferromagnetic magnons.
@@
Figure 1: The models of two antiferromagnetic systems are depicted. The red and blue sites represent the spin-up and spin-down magnetic sublattices of the Néel order. The green atom is non-magnetic and it is placed strictly in the plane with the red and blue sites. The systems include a nearest-neighbor Heisenberg exchange interaction (HEI, $J>0$) and Dzyaloshinskii-Moriya interaction (DMI). Further-neighbor HEI and DMI are indicated by dashed lines. In a simplifying assumption, we exclude any interactions across the green atom. The arrows on the bonds denote the sign of the DMI. The left model possesses a symmetry connecting the two magnetic sublattices, whereas this symmetry is absent in the right model.
II Model
@@
We study two insulating antiferromagnetic systems with the Néel order on the honeycomb-like lattices shown in Fig. (1). These models were studied in Zyuzin2025 for the situation when the systems are metallic. The red and blue sites correspond to ${\bf S}_{1/2}=\pm{\bf S}$ magnetic sublattices. The green atom is non-magnetic, and its role is to alter the first-nearest neighbor Heisenberg (HEI) and Dzyaloshinskii-Moriya (DMI) interactions in its vicinity through the spin-orbit interaction. The green atom is in the plane with the red and blue sites. We also assume that there is no second-nearest neighbor HEI or DMI through the green atom. This is a simplification; however, our results below remain unchanged even if these interactions are non-zero, provided their magnitude differs from that of the corresponding interactions on bonds without the green atom. Exchange Hamiltonian of an antiferromagnet model shown in Fig. (1) is
$H$ $=J\sum_{\langle ij\rangle}S_{i}^{z}S_{j}^{z}+J\sum_{\langle ij\rangle}\cos(\theta_{ij})\left(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}\right)$
$+J\sum_{\langle ij\rangle}\nu_{ij}\sin(\theta_{ij})\left(S_{i}^{x}S_{j}^{y}-S_{i}^{y}S_{j}^{x}\right)$
@@
where $J>0$ and $D\equiv J\sin(\theta)$ is the HEI and DMI between the first neighbors labeled by the $\langle..\rangle$ in the sum. In Fig. (1) $\nu_{ij}=\pm 1$ for the bond along the arrow. If the arrow on the $\langle ij\rangle$ bond is reversed, then $\nu_{ij}=-\nu_{ji}$. If there is no arrow on the bond then $\theta_{ij}=0$. The signs are chosen in correspondence with the underlying spin-orbit coupling. In addition we have added $J^{\prime}$ and $D^{\prime}$, which is the HEI and DMI correspondingly between the second-nearest neighbors, labeled by the $\langle\langle..\rangle\rangle$. We assume that $J^{\prime}<0$, i.e. it is of the ferromagnetic sign. Such a sign supports the Néel order. In addition to this, second-nearest neighbor interactions may be bond and sublattice dependent. The details of these interactions will be specified below for two different models shown in Fig. (1). We stress that there is no external magnetic field in the system.
@@
Figure 3: Spectrum defined in Eq. (15) of the two antiferromagnetic magnon modes, plotted for $\theta=\zeta=0.$ and $t=0.4$ values. Left/right plots correspond to the model I/II. The plots are shown to emphasize the role of the anisotropic exchange interaction given by $t$ in splitting of the magnon modes.
III Magnon thermal Hall effect
The thermal Hall effect of antiferromagnetic magnons has been derived in MatsumotoShindouMurakamiPRB2014 ( also see comment ). Magnon thermal Hall effect is defined as the heat current Luttinger1964 carried by the magnons, which flows in the transverse direction to the applied temperature gradient $\nabla_{\beta}\chi$, namely,
$j^{\mathrm{TH}}_{\alpha}=\frac{1}{2TV}\sum_{{\bf k};n}(\hat{\Omega}_{\alpha\beta;{\bf k}})_{nn}c_{2}\left[(\hat{\sigma}_{3}\hat{E}_{\bf k})_{nn}\right]\nabla_{\beta}\chi,$ (16)
@@ III Magnon thermal Hall effect
$j^{\mathrm{TH}}_{\alpha}=\frac{1}{TV}\sum_{{\bf k};n}\Omega_{\alpha\beta;{\bf k}}\left[c_{2}\left(\delta_{\bf k}+\varepsilon_{\bf k}\right)+c_{2}\left(-\delta_{\bf k}+\varepsilon_{\bf k}\right)\right]\nabla_{\beta}\chi.$ (20)
From ZyuzinKovalevPRL2016 it is known that odd in momentum Berry curvature of antiferromagnetic magnons may be due to the lattice structure (honeycomb lattice), while from Pires (checkerboard lattice with d-wave DMI and finite $\delta_{\bf k}$) it is known that even in momentum Berry curvature may be due to the even in momentum DMI between the sublattices. However, both models, ZyuzinKovalevPRL2016 and Pires don’t allow for non-zero magnon thermal Hall effect because of the symmetry between the magnetic sublattices. These are genuine antiferromagnets and we are going to go over an example of such antiferromagnet in Section IV. Then, in Mook ; CommentAlter an antiferromagent on the rutile lattice was shown to have finite THE due to the even in momentum Berry curvature, similarly to Pires , originating from DMI. It appears that the antiferromagnet on a ruile lattice belongs to the symmetry class of Dzyaloshinskii’s weak ferromagnets Dzyaloshinskii1958 , hence THE of magnons is expected. We will discuss THE of magnons in a collinear weak ferromagnet phase of our model in the Section VI. In the Section V) we show that there is another possibility for non-zero magnon thermal Hall in collinear antiferromagnets shown in Fig. (1) with odd in momentum magnon Berry curvature due to nearest neighbor HEI, and with the even in momentum Berry curvature due to the nearest neighbor DMI. We will show that magnon THE will be non-zero only when all the symmetries between the two magnetic sublattices are broken, and hence the system is a ferrimagnet in this case. Let us discuss the details of models I and II.
IV Mirror-symmetric genuine antiferromagnet
@@
It is clear that the integration over the momentum will make the THE to vanish in this model. It is instructive to see how the Dzyaloshinskii-Turov symmetry Dzyaloshinskii1958 ; Turov1965 arguments support our result. We introduce a Néel vector ${\bf L}={\bf S}_{1}-{\bf S}_{2}$ and the magnetization $\bf M$}. The magnetization $\bf M$} may or may not be related to ${\bf S}_{1}+{\bf S}_{2}$. In the former case the magnetization is due to the canting of the Néel order, while in the latter case the magnetization is carried by magnons (or fermions if the system is metallic). For example, we can have ${\bf M}\neq 0$ even when ${\bf S}_{1}+{\bf S}_{2}=0$. In this paper we are interested in exactly this scenario. We then analyze whether a Dzyaloshinskii’s invariant of the $M_{\alpha}L_{\beta}$ (where $\alpha,\beta=x,y,z$) form can be realized in the system. For it to be non-zero, it has to be invariant under all symmetries of the system. If it is non-zero, it would mean that $L_{\alpha}$ can generate a finite magnetic moment $M_{\beta}$ in the system. In two dimensions we are expecting only $M_{z}$ to be generated. Both vectors transform as pseudovectors, and in addition, $\bf L$} changes sign if the magnetic sublattices are exchanged. For $L_{y}$ and $L_{z}$, a reflection in the yzy-z plane which crosses the vertical bond in the center is the symmetry connecting the two magnetic sublattices. Under this symmetry operation there is no Dzyaloshinskii’s invariant of the $M_{z}L_{z}$ or $M_{z}L_{y}$ form. Indeed, $L_{z}$ and $L_{y}$ are invariant, while $M_{z}$ changes sign. It might appear that a symmetry of $\pi$ rotation about the center of the hexagon and a reflection in the xzx-z plane which crosses the center of the hexagon and time-reversal allows for the $M_{z}L_{x}$ from the list of Dzyaloshinskii’s invariants However, a reflection in the xyx-y plane and time-reversal operation eliminates the $M_{z}L_{x/y}$ combinations. All in all, the system in the left of Fig. (1) is a genuine antiferromagnet with zero magnetic moment, and magnon THE is not expected in this system.
V Ferrimagnet